# Beyond Any-Shot Adaptation: Predicting Optimization Outcome for Robustness Gains without Extra Pay

### v3

@@ F.1 MAML
When it comes to reinforcement learning scenarios, $\mathcal{D}_{\tau}$ corresponds to episodic returns collected from MDPs with either the meta policy or the fast adapted policy. To ensure enough coverage of task space, we adopt a mixture strategy of MPTS and random sampling as an empirical regularizer in all RL scenarios, which is similar to work 33.
@@
Remember that MDP distribution $p(\tau)$ is mostly induced by physical parameters, e.g., mass, gravity, friction, etc., or the reward functions. In each training iteration, the machine learner resamples a batch of MDPs and gets the shared policy to interact with them to collect episodes. Consequently, the query dataset contains the episodes collected with no support dataset. Overall, the policy optimization follows the standard TD3 algorithm 45 due to its sample efficiency and stability.
@@ F.3 Multi-Modal Prompt Learning
Multi-modal prompt learning is based on the backbone of the prompt tuning method MaPLe 44, which we use on both few-shot and many-shot classifications.
@@
The few-shot prompt learning refers to the common few-shot classification setting 28, 85. We integrate the MaPLe backbone with the ProtoNet 85 to fully utilize the support sets in few-shot learning. As illustrated in Section E, we generate the model prediction using both the textual classifiers from the CLIP textual encoder and the visual classifiers from the support set. By freezing the CLIP model parameters, only the prompts are optimized during meta-training. In mathematics, the optimization objective can be formulated as:
@@
Many-shot classification.
The many-shot prompt learning setting refers to the 16-shot classification task proposed in work 96, 97. Based on the MaPLe backbone 44, we again freeze the CLIP model parameters and tune multi-modal prompts. The prompts are optimized on a selected ImageNet subset, with 16 samples from each category.
Note that in the many-shot setting, we do not have pre-defined N-way K-shot tasks, either the splits of support and query sets in each task. Therefore, we replace the “tasks” in the meta-learning setting with training samples. Model predictive task sampling is then achieved through data sampling. In mathematics, the objective can be formulated as:
@@ Appendix G Experimental Setups & Implementation Details
Widely recognized in reinforcement learning is the high sample complexity in policy evaluation, which demands massive interactions with environments, while policy optimization over the MDP distribution makes this even more severe. In N-way K-shot image classification, we can create K-shot classification task from an arbitrary combination of $N$ classes; then the task space complexity $\mathcal{O}\Big{(}C_{M}^{N}\Big{)}$ grows with the number of categories $M$ in image datasets. Meanwhile, challenges arise when gradient updates of foundation models consume substantial computational power and memory with a large batch size. Similar circumstance also occurs in robust finetuning foundation models. We refer the reader to Supplementary Notes Section F/G for all details of experimental setups and implementations.
Table 4: A Summary of the Considered Benchmarks. Here, we list the primary expensive part in task episodic learning for each scenario together with backbone methods. Also note that N-way K-shot image classification and PEFT requires implicit task identifiers while others can directly access explicit task identifiers as the lower dimensional task representation.
@@
      domain randomization       zero-shot       TD3       interactions
      PEFT       many-shot       MaPLe       computation/memory
@@ G.1 Sinusoid Regression
For sinusoid regression, we retain the setup in MAML28, where the few-shot machine learner tries to complete a wave function with the support dataset. In specific, sampling the amplitude $a$ and the phase $b$ configures the wave function, and $10$ data points are uniformly sampled from the interval $[-5.0,5.0]$ coupled with $y=a\sin(x-b)$ to obtain the support dataset. This formulates the $10-\texttt{shot}$ sinusoid regression task.
@@
The machine learner is a neural network with 2 hidden layers of size $40$ with two nonlinear activation units. The task batch for ERM and GDRM is $16$, and that for DRM is $32$ as default. The temperature parameter in GDRM is $\eta=0.001$. The learning rates for the inner loop and the outer loop are $0.001$. The following is about extra optimization details or setups in MPTS. The task identifier’s dimension is $2$ with the latent variable is $\bm{z}\in\mathbb{R}^{x}$. The batch size of the identifier in training is $32$, the Lagrange multiplier is set as $1$, and we use the Adam optimizer with the learning rate $43e-$ 4 to update the risk learner for 20000 step. In sinusoid regression and Meta-RL, we use the standard repository in MAML 28.
@@ G.2 N-way K-shot Image Classification
This is a commonly seen benchmark in few-shot learning. It learns a model that can classify images from $N$ distinct classes with support of $K$ labeled examples for each class. The support dataset as reference is in the form $\mathcal{D}_{\tau}^{S}=\{\{[\bm{x}_{i,k},y_{i,k}=i]\}_{k=1}^{K}\}_{i=1}^{N}$. And the query dataset corresponds to the image information for the model to classify. Hence, for a large image dataset with $M$ classes, the complexity of the task space is $\mathcal{O}\Big{(}C_{M}^{N}\Big{)}$. Here, we include ImageNet-CG 47, ImageNet-CI 47, ImageNet-CS 47, ImageNet-A 48, ImageNet-S 49 and ImageNet-R 50 as the dataset in evaluation.
@@
The machine learner utilizes a prompt learning backbone following MaPLe, with the frozen CLIP model. The task batch for ERM and GDRM is 4, and that for DRM is 8 as default. The temperature parameter in GDRM is 0.001. The learning rate for the outer loop is 0.01. The learning rate for the inner loop follows that in MaPLe 44. The following is about extra optimization details or setups in MPTS. The task identifier is generated by the frozen CLIP text encoder using the input class names, with a dimensionality of 512. The batch size of the identifier in training is 8, the Lagrange multiplier is set as $\beta$, and we use the Adam optimizer with the learning rate 0.01 to update the risk learner.
@@ G.3 Meta-RL
We construct MDP distributions based on Mujoco physics engines 98. These include HalfCheetahVel, HalfCheetahMassVel, Walker2dVel, Walker2dMassVel, and ReacherPos. The HalfCheetahVel and Walker2dVel tasks involve training the cheetah or walker robot to achieve a target velocity. These tasks define the reward function as the negative absolute difference between the robot’s current velocity and the target velocity, supplemented by a control penalty and an alive bonus to facilitate the learning process. The goals and rewards of HalfCheetahMassVel and Walker2dMassVel are the same as those of the corresponding velocity-related tasks, with the additional identifier of varying mass for the cheetah or walker robot. The ReacherPos task tries to move a two-jointed robot arm’s end effector close to a target position, and its reward function is defined as the negative $L$-1 distance between the robot arm’s position and the target position, supplemented by a control cost to ensure robustness.
@@ G.4 Robotic DR
We conduct experiments on LunarLander-v2 and ErgoReacher-v0 environments 9. LunarLander is a 2 degrees of freedom (DoF) environment in which the agent has to land a spacecraft on a designated landing pad without crashing, implemented using Box2D 99. The reward function of LunarLander awards positive rewards for successful landings, negative rewards for crashes, and additional penalties for fuel consumption and deviation from the landing pad, encouraging efficient and controlled landings. ErgoReacher is a 4 DoF arm environment from Golemo et al. 100 in which the arm has to touch a goal with its end effector, implemented in the Bullet Physics Engine 101. The reward function of ErgoReacher includes the negative distance between the end effector’s position and the target, along with other control costs to promote efficient and safe movements. In LunarLander, we randomize the engine strength, while in ErgoReacher, we randomize the joint damping and maximum torque for each of the 4 joints, resulting in a total of 8 parameters. The detailed ranges of the randomized parameters for each environment are provided in Table 3.
@@
G.5 Prompt-Tuning Multimodal Foundation Models as Many-shot Learning
@@
The machine learner follows the prompt learning method MaPLe 44 based on the frozen CLIP model (ViT/B-16). The task batch for ERM and GDRM is 4, and that for DRM is 8 as default. The temperature parameter in GDRM is 0.001. The learning rate for the outer loop is 0.005. The learning rate for the inner loop follows that in MaPLe 44. The following is about extra optimization details or setups in MPTS. As for the neural architecture of the risk learner, the encoder is a neural network with 5 hidden layers with 4 ReLU nonlinear activation units, and the decoder is a neural network with 4 hidden layers with 3 nonlinear activation units. The task identifier’s dimension is 512 with the latent embedding from CLIP encoders. The batch size of the identifier in training is 8, the Lagrange multiplier is set as $\beta$, and we use the Adam optimizer with the learning rate 0.005 to update the risk learner for 8000 steps. During prompt-tuning, we set the task batch for ERM and DRM as $\mathcal{B}=4$ (For implementation simplicity, the data loader samples $8$ tasks and then randomly keeps half without ranking to optimize). The task batch for DRM is $\mathcal{B}=8$ before the filtering operation; DRM filters half to optimization. Similarly, that for MPTS is $\mathcal{B}=8$ and only $4$ tasks are screened to optimize.
@@
      ImageNet-CG 47       $\text{CVaR}_{0.9}$       77.02       77.76       77.61       78.04       +1.02
@@
      ImageNet-CI 47       $\text{CVaR}_{0.9}$       80.24       80.47       80.15       80.97       +0.73
@@
      ImageNet-CS 47       $\text{CVaR}_{0.9}$       76.63       77.71       76.18       78.4       +1.77
@@
      ImageNet-A 48       $\text{CVaR}_{0.9}$       76.69       77.39       77.40       77.95       +1.26
@@
      ImageNet-S 49       $\text{CVaR}_{0.9}$       82.63       83.67       83.11       84.82       +2.21
@@
      ImageNet-R 50       $\text{CVaR}_{0.9}$       88.31       88.84       87.71       89.79       +1.48
@@
     ImageNet 51 (ID)      $\text{CVaR}_{0.9}$      31.68      32.46      31.38      32.5      +0.82
@@
     ImageNet-A 48 (OOD)      $\text{CVaR}_{0.9}$      15.33      15.46      15.6      18.44      +3.11
@@
     ImageNet-R 50 (OOD)      $\text{CVaR}_{0.9}$      26      28.06      25.9      28.19      +2.19
@@
     ImageNet-S 49 (OOD)      $\text{CVaR}_{0.9}$      12.24      13.16      12.22      13.63      +1.39
@@
     ImageNet-V 52 (OOD)      $\text{CVaR}_{0.9}$      24.6      25.5      24.2      25.90      +1.3
@@ Appendix I Competing Interests & Author Contributions
Under the supervision of professor X.J., Q.C.W. conceptualized the idea of MPTS, designed the computational framework, formulated the mathematical part, and wrote the draft. Z.X. and J.S. implemented MPTS in sinusoid regression, prompt-based few-shot image classification, and PEFT image classification, collected experimental results to visualize, and added implementation details in Supplementary Material. Y.M., Y.Q., and Y.L. implemented MPTS in sinusoid regression, Meta-RL and robotic DR, collected experimental results to visualize, and added implementation details in Supplementary Material. X.J. supervised the progress of MPTS, organized technical discussions, reviewed and revised the original draft. All authors have read the manuscript and approved the public version.

### v5

@@ F.1 MAML
When it comes to reinforcement learning scenarios, $\mathcal{D}_{\tau}$ corresponds to episodic returns collected from MDPs with either the meta policy or the fast adapted policy. To ensure enough coverage of task space, we adopt a mixture strategy of MPTS and random sampling as an empirical regularizer in all RL scenarios, which is similar to work 17.
@@
Remember that MDP distribution $p(\tau)$ is mostly induced by physical parameters, e.g., mass, gravity, friction, etc., or the reward functions. In each training iteration, the machine learner resamples a batch of MDPs and gets the shared policy to interact with them to collect episodes. Consequently, the query dataset contains the episodes collected with no support dataset. Overall, policy optimization follows the standard TD3 algorithm 37 due to its sample efficiency and stability.
@@ F.3 Multi-Modal Prompt Learning
Multi-modal prompt learning is based on the backbone of the prompt tuning method MaPLe 36, which we use on both few-shot and SFT for image classifications.
@@
The few-shot prompt learning refers to the common few-shot classification setting 29, 83. We integrate the MaPLe backbone with the ProtoNet 83 to fully utilize the support sets in few-shot learning. As illustrated in Section E, we generate the model prediction using both the textual classifiers from the CLIP textual encoder and the visual classifiers from the support set. By freezing the CLIP model parameters, only the prompts are optimized during meta-training. In mathematics, the optimization objective can be formulated as:
@@
SFT for image classification.
The prompt learning setting refers to the 16-shot classification task proposed in work 94, 95. Based on the MaPLe backbone 36, we again freeze the CLIP model parameters and tune multi-modal prompts. The prompts are optimized on a selected ImageNet subset, with 16 samples from each category.
Note that in the SFT setting, we do not have pre-defined N-way K-shot tasks, either the splits of support and query sets in each task. Therefore, we replace the “tasks” in the meta-learning setting with training samples. Model predictive task sampling is then achieved through data sampling. In mathematics, the objective can be formulated as:
@@ Appendix G Experimental Setups & Implementation Details
Widely recognized in reinforcement learning is the high sample complexity in policy evaluation, which demands massive interactions with environments, while policy optimization over the MDP distribution makes this even more severe. In N-way K-shot image classification, we can create K-shot classification task from an arbitrary combination of $N$ classes; then the task space complexity $\mathcal{O}\Big{(}C_{M}^{N}\Big{)}$ grows with the number of categories $M$ in image datasets. Meanwhile, challenges arise when gradient updates of foundation models consume substantial computational power and memory with a large batch size. Similar circumstance also occurs in robust finetuning foundation models.
Table 3: A Summary of the Considered Benchmarks. Here, we list the primary expensive part in task episodic learning for each scenario together with backbone methods. Also note that N-way K-shot image classification and SFT requires implicit task identifiers while others can directly access explicit task identifiers as the lower dimensional task representation.
@@
      Robotic DR       zero-shot       TD3       interactions
      SFT       many-shot       MaPLe       computation/memory
@@ G.1 Sinusoid Regression
For sinusoid regression, we retain the setup in MAML29, where the few-shot machine learner tries to complete a wave function with the support dataset. In specific, sampling the amplitude $a$ and the phase $b$ configures the wave function, and $10$ data points are uniformly sampled from the interval $[-5.0,5.0]$ coupled with $y=a\sin(x-b)$ to obtain the support dataset. This formulates the $10-\texttt{shot}$ sinusoid regression task.
@@
The machine learner is a neural network with 2 hidden layers of size $40$ with two nonlinear activation units. The task batch for ERM and GDRM is $16$, and that for DRM is $32$ as default. The temperature parameter in GDRM is $\eta=0.001$. The learning rates for the inner loop and the outer loop are $0.001$. The following is about extra optimization details or setups in MPTS. The task identifier’s dimension is $2$ with the latent variable is $\bm{z}\in\mathbb{R}^{x}$. The batch size of the identifier in training is $32$, the Lagrange multiplier is set as $1$, and we use the Adam optimizer with the learning rate $43e-$ 4 to update the risk learner for 20000 step. In sinusoid regression and Meta-RL, we use the standard repository in MAML 29.
@@ G.2 N-way K-shot Image Classification
This is a commonly seen benchmark in few-shot learning. It learns a model that can classify images from $N$ distinct classes with support of $K$ labeled examples for each class. The support dataset as reference is in the form $\mathcal{D}_{\tau}^{S}=\{\{[\bm{x}_{i,k},y_{i,k}=i]\}_{k=1}^{K}\}_{i=1}^{N}$. And the query dataset corresponds to the image information for the model to classify. Hence, for a large image dataset with $M$ classes, the complexity of the task space is $\mathcal{O}\Big{(}C_{M}^{N}\Big{)}$. Here, we include ImageNet-CG 44, ImageNet-CI 44, ImageNet-CS 44, ImageNet-A 45, ImageNet-S 27 and ImageNet-R 46 as the dataset in evaluation.
@@
Explicit $\bm{\tau}$ are unavailable to specify the task; however, it can be approximately resolved by describing the identifier through a small reference model. Specifically, we leverage CLIP’s text encoder to obtain $\bm{\tau}\approx[\text{CLIP}_{\text{text}}(\mathcal{C}_{1}),\dots,\text{CLIP}_{\text{text}}(\mathcal{C}_{K})]$ with the tokenizations of $K$ class texts $\mathcal{C}_{1:K}$ from $\mathcal{D}_{\tau}^{S}$. The machine learner utilizes a prompt learning backbone following MaPLe, with the frozen CLIP model. The task batch for ERM and GDRM is 4, and that for DRM is 8 as default. The temperature parameter in GDRM is 0.001. The learning rate for the outer loop is 0.01. The learning rate for the inner loop follows that in MaPLe 36. The following is about extra optimization details or setups in MPTS. The task identifier is generated by the frozen CLIP text encoder using the input class names, with a dimensionality of 512. The batch size of the identifier in training is 8, the Lagrange multiplier is set as $\beta$, and we use the Adam optimizer with the learning rate 0.01 to update the risk learner.
@@ G.3 Meta-RL
We construct MDP distributions based on Mujoco physics engines 96. These include HalfCheetahVel, HalfCheetahMassVel, Walker2dVel, Walker2dMassVel, and ReacherPos. The HalfCheetahVel and Walker2dVel tasks involve training the cheetah or walker robot to achieve a target velocity. These tasks define the reward function as the negative absolute difference between the robot’s current velocity and the target velocity, supplemented by a control penalty and an alive bonus to facilitate the learning process. The goals and rewards of HalfCheetahMassVel and Walker2dMassVel are the same as those of the corresponding velocity-related tasks, with the additional identifier of varying mass for the cheetah or walker robot. The ReacherPos task tries to move a two-jointed robot arm’s end effector close to a target position, and its reward function is defined as the negative $L$-1 distance between the robot arm’s position and the target position, supplemented by a control cost to ensure robustness.
@@ G.4 Robotic DR
We conduct experiments on LunarLander-v2 and ErgoReacher-v0 environments 35. LunarLander is a 2 degrees of freedom (DoF) environment in which the agent has to land a spacecraft on a designated landing pad without crashing, implemented using Box2D 97. The reward function of LunarLander awards positive rewards for successful landings, negative rewards for crashes, and additional penalties for fuel consumption and deviation from the landing pad, encouraging efficient and controlled landings. ErgoReacher is a 4 DoF arm environment from 98 in which the arm has to touch a goal with its end effector, implemented in the Bullet Physics Engine 99. The reward function of ErgoReacher includes the negative distance between the end effector’s position and the target, along with other control costs to promote efficient and safe movements. In LunarLander, we randomize the engine strength, while in ErgoReacher, we randomize the joint damping and maximum torque for each of the 4 joints, resulting in a total of 8 parameters. The detailed ranges of the randomized parameters for each environment are provided in Table 2.
@@
G.5 Prompt-Tuning Multimodal Foundation Models
@@
The machine learner follows the prompt learning method MaPLe 36 based on the frozen CLIP model (ViT/B-16). The task batch for ERM and GDRM is 4, and that for DRM is 8 as default. The temperature parameter in GDRM is 0.001. The learning rate for the outer loop is 0.005. The learning rate for the inner loop follows that in MaPLe 36. The following is about extra optimization details or setups in MPTS. As for the neural architecture of the risk learner, the encoder is a neural network with 5 hidden layers with 4 ReLU nonlinear activation units, and the decoder is a neural network with 4 hidden layers with 3 nonlinear activation units. The task identifier’s dimension is 512 with the latent embedding from CLIP encoders. The batch size of the identifier in training is 8, the Lagrange multiplier is set as $\beta$, and we use the Adam optimizer with the learning rate 0.005 to update the risk learner for 8000 steps. During prompt-tuning, we set the task batch for ERM and DRM as $\mathcal{B}=4$ (For implementation simplicity, the data loader samples $8$ tasks and then randomly keeps half without ranking to optimize). The task batch for DRM is $\mathcal{B}=8$ before the filtering operation; DRM filters half to optimization. Similarly, that for MPTS is $\mathcal{B}=8$ and only $4$ tasks are screened to optimize.
@@
      ImageNet-CG 44       $\text{CVaR}_{0.9}$       77.02       77.76       77.61       78.04       +1.02
@@
      ImageNet-CI 44       $\text{CVaR}_{0.9}$       80.24       80.47       80.15       80.97       +0.73
@@
      ImageNet-CS 44       $\text{CVaR}_{0.9}$       76.63       77.71       76.18       78.4       +1.77
@@
      ImageNet-A 45       $\text{CVaR}_{0.9}$       76.69       77.39       77.40       77.95       +1.26
@@
      ImageNet-S 27       $\text{CVaR}_{0.9}$       82.63       83.67       83.11       84.82       +2.21
@@
      ImageNet-R 46       $\text{CVaR}_{0.9}$       88.31       88.84       87.71       89.79       +1.48
@@
     ImageNet 26 (ID)      $\text{CVaR}_{0.9}$      31.68      32.46      31.38      32.5      +0.82
@@
     ImageNet-A 45 (OOD)      $\text{CVaR}_{0.9}$      15.33      15.46      15.6      18.44      +3.11
@@
     ImageNet-R 46 (OOD)      $\text{CVaR}_{0.9}$      26      28.06      25.9      28.19      +2.19
@@
     ImageNet-S 27 (OOD)      $\text{CVaR}_{0.9}$      12.24      13.16      12.22      13.63      +1.39
@@
     ImageNet-V 47 (OOD)      $\text{CVaR}_{0.9}$      24.6      25.5      24.2      25.90      +1.3
@@ Appendix I Competing Interests & Author Contributions
Under the supervision of professor X.J., Q.C.W. conceptualized the idea of MPTS, designed the computational framework, formulated the mathematical part, and wrote the draft. Z.X. and J.S. implemented MPTS in sinusoid regression, prompt-based few-shot image classification, and SFT image classification, collected experimental results to visualize, and added implementation details in Supplementary Material. Y.M., Y.Q., and Y.L. implemented MPTS in sinusoid regression, Meta-RL and robotic DR, collected experimental results to visualize, and added implementation details in Supplementary Material. X.J. supervised the progress of MPTS, organized technical discussions, reviewed and revised the original draft. All authors have read the manuscript and approved the public version.